OpenStudy (anonymous):

1/3 log 4z+2 log 4(2z+1)

OpenStudy (anonymous):

\[1/3 \log_{4} z+2 \log_{4}(2z+1) \]

OpenStudy (anonymous):

\[18\log_{\sqrt{x}} +9\log_{\sqrt[3]{x}}-\log 10\]